OpenStudy (anonymous):

solve the equation x(4x-3)=7

OpenStudy (anonymous):

therefore x =7 or 4x-3 = 7 4x=10 x=2.5

OpenStudy (uri):

Mutliply the whole bracket with x x*4 and -x*3 what do you egt? @Roadhoue222

OpenStudy (anonymous):

i know my 1st part of the answer is -1, i can't get the 2nd part, and no decimals

OpenStudy (uri):

get*

OpenStudy (uri):

DO what i said..

OpenStudy (anonymous):

please explain uri

OpenStudy (uri):

What is \[x \times 4 -x \times 3?\]

OpenStudy (anonymous):

4x^2-3x

OpenStudy (uri):

@Roadhoue222 Are you sure this is the question? :)

OpenStudy (anonymous):

yes @uri the question x(4x-3)=7

OpenStudy (anonymous):

solve the equation

OpenStudy (uri):

Yes it's 4x sorry i made a typo you're right! Now 4x^2-3x=7 Arrange the equation 4x^2-3x-7=0

OpenStudy (uri):

Well we can't factorize it..do you know Quad formula?

OpenStudy (anonymous):

is it (x+1)(3x-7)

OpenStudy (anonymous):

no i don't